### v1

# Unrolled Creative Adversarial Network For Generating Novel Musical Pieces

Abstract
Music generation has been established as a prominent topic in artificial intelligence and machine learning [2, 13] over recent years. In most recent works [4, 7, 3], RNN-based neural network methods have been applied for sequence generation. In contrast, generative adversarial networks (GANs) and their counterparts have been explored by very few researchers [10] for music generation.
In this paper, a classical system was employed alongside a new system to generate creative music. Both systems were designed based on adversarial networks to generate music by learning from examples. The classical system was trained to learn a set of music pieces without differentiating between classes, whereas the new system was trained to learn the different composers and their styles to generate a creative music piece by deviating from the learned composers’ styles.
The base structure utilized was generative adversarial networks (GANs), which are capable of generating novel outputs given a set of inputs to learn from and mimic their distribution. It has been shown in previous work that GANs are limited in their original design with respect to creative outputs [5]. Building on the Creative Adversarial Networks (CAN) [5], this work applied them in the music domain rather than the visual art domain. Additionally, unrolled CAN was introduced to prevent mode collapse. Experiments were conducted on both GAN and CAN for generating music, and their capabilities were measured in terms of deviation from the input set.
1 Introduction
Machines have always been shown as devices capable of replicating certain learned behaviours. It has been a dream to make creative machines and is of much desire in the machine learning scope and artificial intelligence. After all if we want machines to mimic human behaviour, one the most important human traits is creativity. Take driving for example; humans are capable of using their creativity and creative problem solving skills, to avoid unexpected road events, preventing accidents. That is a skill that all self driving car developers hope to accomplish. The art of generating music is a very complex problem [10] and is a different challenge than that of text generation. It includes finding a rhythm between the musical notes and learning polyphony and many more challenges.
To study style in music lets take a look at what style means in language. When one speaks or writes their style is defined by their selection of words an vocabulary, their choice of grammar and linguistics, all governed within constraints of the language and dialect one has learned, and have not created. However, since constraints can allow for interpretations of different realizations it produces variety. Therefore, a pattern in a style of writing or a piece of music may not relate in every aspect, only some aspects are selected to convey pattern replication. 1,2 shows two examples of musical pieces replicating a pattern of using pitches (8-7/2-1) typically preferred and chosen by classical musicians [9].
Figure 1: Symphony No.46 in B major Haydn
Figure 2: Mozart Piano Quartet in G minor
There are two clefs in a music sheets (3) that are used to differentiate between high and bass notes and are called the treble clef and the bass clef. The treble clef is usually played with the right hand on the piano and the bass clef is usually played with the left hand on the piano.
Figure 3: Treble clef and Bass clef
In music sheets some note may be represented different graphically yet representing the same note. And sometimes certain music sheets may look identical graphically while actually are representing different notes. That is due to the change in the clef (Check figure 4).
A professional musician must be able to identify the similarities and differences in the notes. That comes naturally since professional musicians are regularly reading sheet music.
Figure 4: Same-Different exercise example [12]
2 Methodology
In this section the methodology of music GAN, CAN and unrolled CAN pipeline is discussed. Starting with the data sets used, data preprocessing technique, and the model architecture.
2.1 Data set
Three data sets were used in this project; jazz midi data set, classical music midi data set classified by composers, and Arabic music midi data set. All the data sets are included in the following github repository. Both the jazz data set and the classical music data set were retrieved from kaggle. The Arabic data set was collected manually and cleaned to cancel out some of the repetitive notes such as drums.
2.2 Data Preprocessing
The data used in this project contain midi files. In order to pass them to the GAN or the CAN the midi files were converted to images by converting each note to a pixel location. The resulting image file can be converted back to midi file without losing any of the information. (see figure 5)
Figure 5: Sample of musical piece by the composer Albeniz in image format
2.3 Music GAN architecture
The architecture used for the music GAN is presented here. Two networks were included in the architecture: a generator network and a discriminator network. A random vector $z$ was used as input by the generator to produce an image representing a MIDI file. The generated image was then passed to the discriminator, which attempted to distinguish between the real image dataset and the generated (fake) images. The GAN was trained on the three datasets separately, resulting in three distinct generators: one capable of producing jazz musical pieces, another capable of generating classical music pieces, and the third trained on manually collected and cleaned data to generate Arabic musical pieces.
Figure 6: GAN architecture
A detailed architecture of Discriminator and Generator. 7,8
Figure 7: GAN Discriminator block
Figure 8: GAN Generator block
2.4 Music CAN and unrolled CAN architecture
In the music CAN the architecture[11] [5]is still made up of two neural networks that compete against each other. However, the loss here stresses on the generator motivating it to produce a musical piece that cannot be classified as belonging to one class. In other words, the discriminator should be capable of classifying the original real data to their composers or class label. The discriminator should not be able to find a suitable class for the generated piece of music. Thus the input data must be labeled by the class it belongs to in order to reduce the loss on real data classification. The generator architecture is similar to that of the GAN where it takes a random z vector as an input and generates an image representing the midi file.
Figure 9: CAN architecture
A detailed architecture of Discriminator and Generator. 10, 11
Figure 10: CAN Discriminator block
Figure 11: CAN Generator block
For the CAN generator architecture, a size 100 random $z$ vector is converted into a $512\times 4\times 4$ tensor through the use of convolution and batch normalization layers. Convolution and batch normalization are then applied sequentially to produce a $256\times 8\times 8$ tensor, followed by a $128\times 32\times 32$ tensor, and subsequently a $64\times 64\times 64$ tensor. Finally, convolution and batch normalization are applied once more to generate a $1\times 128\times 128$ tensor, which represents the graphical form of the MIDI file.
For the discriminator, two convolution layers are employed with batch normalization applied to both, followed by a leaky ReLU activation and a fully connected layer.
3 Technical Details
3.1 Generative Adversarial Networks
GAN was first present by Ian J. Goodfellow on 2014 where two deep neural networks play a minmax game over the value function $V(G,D)$ [6]. Where G is the Generator and D is the Discriminator. And z is a random vector supplied to the generator as input.
[…]
$\mathbb{E}_{z\sim p_{z}(z)}[log(1-D(G(z)))]$ (1)
3.2 Creative Adversarial Networks
CAN was presented recently where a modification to the value function to include cross entropy loss over the classes of the generated output ($G(z)$)[5]. Thus now the architecture includes classification and therefore, the loss also includes classification loss. And in order to convey creativity the authors used class ambiguity loss. The class ambiguity loss is achieved by maximizing the class posterior entropy. And hence, the value function is as follows
[…]
$\mathbb{E}_{x,\hat{c}\sim p_{data}}[logD_{r}(x)+logD_{c}(c=\hat{c}|x)]+$
$\mathbb{E}_{z\sim p_{z}}[log(1-D_{r}(G(z)))-\sum_{k=1}^{K}(\frac{1}{K}log(D_{c}(c_{k}|G(z)))+$
$(1-\frac{1}{K})log(1-D_{c}(c_{k}|G(z))))]$
Where as above z here is the random vector. And c here represents the class label.
3.3 Unrolling Creative Adversarial Networks
The unrolled CAN is a combination between unrolled GAN and CAN architecture where this structure allows the generator network to look into the future of the discriminator network and optimize over the future discriminator but not updating the discriminator just yet[8]. In other words, the generator is a number (unrolling step numbers may vary) of steps ahead than the discriminator. The result is a reduction in the strength of the discriminator allowing the generator to catch up. Figure 12 shows the result of the regular CAN network after 20 epochs and the results after the modification was introduced making it an unrolled CAN.
Figure 12: Example of CAN, with k=0, output converging to almost identical result after 20 epochs
The way the parameters are updated in GAN is described by the set of equations below
$G\leftarrow G-\eta\frac{dV(D(G),G)}{dG}$ (3)
$D\leftarrow D-\eta\frac{dV(D,G)}{dD}$ (4)
where $V(D,G)$ is the value function. The updates in parameters after unrolling will be as follows
$G\leftarrow G-\eta\frac{dV(D^{k}(G),G)}{dG}$ (5)
$D\leftarrow D-\eta\frac{dV(D,G)}{dD}$ (6)
Where $\eta$ here represents the learning rate. and $k$ is the number of unrolling steps. See figure 13 for the changes due to Unrolled-CAN.
Figure 13: Example of unrolled CAN, with k=1, output converging to varying result after 20 epochs
4 Results
This section will discuss and compare the results of the three architectures; GAN, CAN, and unrolled CAN. The results are presented after post-processing the images and converting them to musical notes extracted from their corresponding midi files.
4.1 Scoring novelty generated musical pieces
In this paper, the expert gate technique [1] was followed, where an auto-encoder was trained on the reconstruction of the real dataset. The expert gate was trained to become proficient in reconstructing the real dataset, which was then used to evaluate the novelty of the generated images. The Mean Squared Error (MSE) was used as the scoring metric. Higher scores indicated greater creativity and deviation from the training set. After the training process, the reconstruction error of the real dataset was observed to have an average of 0.01 MSE. The network architecture is shown in Figure 14.
Figure 14: Auto-encoder network architecture
4.2 Music GAN Results
The music GAN was capable of generating musical pieces that sounded and scored closely to the dataset. The output shown below in Figure 15 represents the result of training on the Jazz dataset. The generated data files exhibited very low reconstruction error, with an average of 0.02 MSE. However, the generated music produced by the GAN architecture was only able to capture the treble clef and was incapable of producing the bass clef.
Figure 15: Output musical note from classical GAN architecture trained on the Jazz dataset
4.3 Music CAN Results
The CAN network was trained on the classical music dataset, where each composer was assigned as a class label. The goal was to generate new classical music that deviates from the styles of all the classical music composers. From the results shown in Figure 16, an improvement over the GAN network can be observed, as the bass clef was captured with greater coherence between the notes. The average score of the generated music set was found to be 0.013 MSE. This indicates that, although the generated music is novel and exhibits high entropy between composers’ styles, it remains relatively close to the original dataset.
Figure 16: Output musical note from CAN architecture trained on the classical music dataset
4.4 Unrolled Music CAN Results
The unrolled CAN architecture combined the strengths of both GAN and CAN, resulting in higher scores while maintaining high entropy between composers’ styles. The average score of the music generated by the unrolled CAN was observed to be 0.043 MSE. This improvement is highlighted in Figure 17.
Figure 17: Output musical note from Unrolled CAN architecture trained on the classical music dataset
4.5 Discussion
It is observed that unrolled CAN is capable of deviating from learned styles and also deviating from the data set while still being able to trick the discriminator network.
Network Max Novelty Score
GAN 0.029
CAN 0.013
Unrolled CAN 0.043
Table 1: Results. Unrolled CAN is better.
Non the less the CAN architecture was very good at generating music that deviates from the composers’ styles but converging to the generation of similar results leading to mode collapse which could be the reason why the average reconstruction error was very low.
4.6 Reproducing the Results
All the codes related to this work have been uploaded to the following GitHub repository: GITHUB REPO.
To train the Unrolled CAN model on the dataset, the file GAN_reconstructed.ipynb located in the folder CAN_models should be executed. Sample data has been provided in the classical_data_midi folder. The files training.py and model.py contain all the model-related configurations. MIDI datasets can be preprocessed into images using the data_pre_processing.ipynb file.
To train the second-best model, the DCGAN, the folder GAN_models should be accessed, and the file GAN-Copy2.ipynb must be run. Before execution, the proper folder path containing the image representation of the MIDI datasets should be specified in the GAN-Copy2.ipynb file. Some generated results for both models have been uploaded to their respective folders.
The novelty_score.ipynb file has been utilized to calculate the novelty score, as discussed in [1]. This file is located in the home directory.
The datasets used in this work are publicly available: Jazz MIDI Dataset and Classical MIDI Dataset.
5 Conclusion
This marks a new artistic era where machines are introduced into the creative domain. Machines are now capable of contributing to the artistic and imaginative sectors, areas where their involvement was previously limited. It is encouraging to witness machines evolving in artificial intelligence and generating ideas that humans might find challenging to perceive. While humans require years to master music theory and play fluently, the algorithm presented in this paper demonstrates the ability to learn musical instruments fluently within minutes and produce music in seconds.
This paper explored different ways musical styles are classified and reviewed multiple architectures. It was demonstrated that the GAN structure [6] and CAN [5] are fully capable of generating music. Additionally, a novel modification to the CAN was introduced by unrolling it, as detailed in [8]. This project combined art and science, leveraging machine learning to create novel and innovative music. Such advancements could inspire new playlists for hotels, aid singers and music composers in seeking new inspirations, and reaffirm the purpose of machines as tools to benefit humanity.
Future improvements could include expanding the dataset to incorporate a wider variety of musical styles and introducing multiple instruments within a single piece. This would enable the development of an unrolled concert CAN capable of orchestrating complex compositions using multiple instruments.